OpenStudy (abbycross167):

Can someone please help me on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

For questions 14, 15 and 16, use f(x) = -3x + 2 14. find f(2)

OpenStudy (abbles):

Plug in 2 for x. f(2) = -3(2) + 2 f(2) = -6 + 2 f(2) = -4

OpenStudy (abbycross167):

Thank you!! so for 15.Find f(0) these would be the steps? f(0) = -3(0) + 2 f(0)= 0+2 f(0)= 2?

OpenStudy (abbycross167):

@Abbles

OpenStudy (abbles):

Perfect!

OpenStudy (abbycross167):

awesome!! thank you so much for making this easy to understand!!

OpenStudy (abbles):

no problemo ;) let me know if you have any questions

OpenStudy (abbycross167):

I have a few more that I need some help with if you don't mind helping me. I'm about to post a new question